OpenStudy (anonymous):

write the partial fraction decomposition of (X^3 + 4x - 1)/(X^3 - X^2)

OpenStudy (anonymous):

hmmm the degrees are the same so your have to divide first

OpenStudy (dumbcow):

yeah i think polynomial division is what you need here

OpenStudy (anonymous):

yikes and it is a drag after that to because of repeated root.

OpenStudy (dumbcow):

can't you then use partial fractions with remainder

myininaya (myininaya):

cow is right

myininaya (myininaya):

you need long division here first

OpenStudy (anonymous):

yes sure. but you will still have a repeated root because denominator is \[x^2(x-1)\] so you are going to have to write \[\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-1}\] have fun. i would use technology.

OpenStudy (anonymous):

ok i did the division and get 1 + (X^2 + 4X -2)/(X^3 - X^2)

OpenStudy (anonymous):

ok so now you set up \[\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-1}=\frac{x^2 + 4x -2}{x^3 - x^2}\] assuming your division is right

myininaya (myininaya):

1 ----------------------- x^3-x^2 | x^3 +0x^2 +4x -1 -(x^3 -x^2) ---------------------------- 1x^2+4x-1 so we have \[\frac{x^3+4x-1}{x^3-x^2}=1+\frac{x^2+4x-1}{x^3-x^2}\] now we can do partial fractions

OpenStudy (anonymous):

@myininaya he got something else. you sure you are right?

OpenStudy (dumbcow):

you can't do \[\frac{A}{x-1} + \frac{Bx+C}{x^{2}}\] ?

OpenStudy (anonymous):

nope

myininaya (myininaya):

am i wrong?

OpenStudy (anonymous):

you have to take all powers.

OpenStudy (anonymous):

oh i don't know because i didn't do it. he wrote something else. give me a second

myininaya (myininaya):

i think i did my division right

OpenStudy (anonymous):

hes write i mistype

OpenStudy (anonymous):

sorry

myininaya (myininaya):

i'm a her

myininaya (myininaya):

lol i think im gonna change my icon to something more girly

OpenStudy (anonymous):

why should i doubt you (honorary male tonight i see)

OpenStudy (anonymous):

nothing too lewd i trust

OpenStudy (anonymous):

\[\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-1}=\frac{x^2 + 4x -1}{x^3 - x^2}\]

OpenStudy (anonymous):

i mistype that too i meant she not hes

myininaya (myininaya):

:)

OpenStudy (anonymous):

\[Ax(x-1)+B(x-1)+Cx^2=x^2+4x-1\] right so far?

OpenStudy (anonymous):

i let x = 0 get B = 1 right off the bat

OpenStudy (dumbcow):

satellite i think either way might work \[= 1 + \frac{-3x+1}{x^{2}} + \frac{4}{x-1}\] it matches same graph, so it works

OpenStudy (anonymous):

sure but you still have that middle term right? so you can do pfd on it as well

OpenStudy (dumbcow):

oh yeah -3/x and 1/x^2

OpenStudy (anonymous):

i guess what i am saying is "you are not done yet"

OpenStudy (anonymous):

ok so i guess we are done right? it is \[1-\frac{3}{x}+\frac{1}{x^2}+\frac{4}{x-1}\]

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

i only got as far as B = 1, but i guess armed with that the other two are easy. for example let x = 1, then c = 4 and we know that A+C=1 so A = -3 and we have them all

OpenStudy (anonymous):

wow! thanks i need more practice